(3i)(2-5i)
= 6i-15i^2
= sqrt-6 - (-1)sqrt15
= sqrt-6 + sqrt15 << This was my answer, however my book has a more sensible answer which is: 15+6i.
Can I get help on what I am doing wrong please?
Thank you for your time.

Answer by Alan3354(69443)

(3i)(2-5i)
= 6i-15i^2
==================
= 6i - 15*(-1)
= 6i + 15
= 15 + 6i
-------------------------
"sensible" is not an issue.
--------------------------------------
= sqrt-6 - (-1)sqrt15
= sqrt-6 + sqrt15 << This was my answer, however my book has a more sensible answer which is: 15+6i.